Arnold Vimka is a venture capitalist facing two alternative investment opportunities. He intends to invest $1 million in a start-up firm. He is nervous, however, about future economic volatility. He asks you to analyze the following financial data for the past year’s operations of the two firms he is considering and give him some business advice.

Company Name
Larson Benson
Variable cost per unit (a) $ 19.00 $ 9.50
Sales revenue (8,600 units × $29.00) $ 249,400 $ 249,400
Variable cost (8,600 units × a) (163,400 ) (81,700 )
Contribution margin $ 86,000 $ 167,700
Fixed cost (24,000 ) (105,700 )
Net income $ 62,000 $ 62,000

Required

  1. Use the contribution margin approach to compute the operating leverage for each firm.

  2. If the economy expands in coming years, Larson and Benson will both enjoy a 12 percent per year increase in sales, assuming that the selling price remains unchanged. Compute the change in net income for each firm in dollar amount and in percentage. (Note: Since the number of units increases, both revenue and variable cost will increase.)

  3. If the economy contracts in coming years, Larson and Benson will both suffer a 12 percent decrease in sales volume, assuming that the selling price remains unchanged. Compute the change in net income for each firm in dollar amount and in percentage. (Note: Since the number of units increases, both revenue and variable cost will increase.)
Answer #1

Solution a:

Operating leverage = Contribution margin / net income

Operating leverage - Larson = $86,000 / $62,000 = 1.387096

Operating leverage- Benson = $167,700 / $62,000 = 2.704839

Solution b:

If sales increases by 12% for larson then net income increases by = % increase in sales * Operating leverage = 12% * 1.387096 = 16.645152%

Dollar change in net income of larson = $62,000 * 16.645152% = $10,320

If sales increases by 12% for Benson then net income increases by = % increase in sales * Operating leverage = 12% * 2.704839 = 32.458068%

Dollar change in net income of Benson = $62,000 * 32.458068% = $20,124

Solution c:

If sales decreases by 12% for larson then net income deccreases by = % decrease in sales * Operating leverage = 12% * 1.387096 = 16.645152%

Decrease in net income of larson = $62,000 * 16.645152% = $10,320

If sales decreases by 12% for Benson then net income decreases by = % decrease in sales * Operating leverage = 12% * 2.704839 = 32.458068%

Decrease in net income of Benson = $62,000 * 32.458068% = $20,124